OpenStudy (sugarmochi):

2 - 4x + 4y and 2y += 3x + 13 please help. I have a test in a few hrs and i still dont understand equations ;-;

OpenStudy (sugarmochi):

It said to solve the following simultaneously

OpenStudy (chaise):

Your first equation is incomplete, there should be an equals sign, we need to know where it is to help solve.

OpenStudy (sugarmochi):

oh oops its 2 - 4x = 4y

OpenStudy (sugarmochi):

and the 2nd one is 2y = 3x + 13

OpenStudy (anonymous):

solve first one for x then put the value of the x in the second equation to find the value of y..... then use any equation to solve for x using the value of y

OpenStudy (sugarmochi):

umm so 4x = 4y + 2?

OpenStudy (sugarmochi):

then, 3( 4y + 2) = 2y

OpenStudy (anonymous):

2 - 4x = 4y and 2y = 3x + 13 Firstly, take the first equation, divide both sides by 4 then you will get : 0.5 - x = y Then plug this into the second equation you will get : 2(0.5-x) = 3x+13 which is : 1 - 2x = 3x + 13 get the -2x from the left hand side to the right hand one you will get : 1 = 5x + 13 the solving for x you will get : x = -12/5 which is -2.4 then substituting back by x = -2.4 into y = 0.5 - x you will get y which will be : y = 0.5 - (-2.4) = 0.5 + 2.4 = 2.9 Let me know if you got it :)

OpenStudy (sugarmochi):

Ahhh i get it now. Thankyou so much, also does this method apply to all of these simultaneous equations?

OpenStudy (anonymous):

Yes just apply these few simple steps carefully and you will get a full mark in all the questions of this type :)

OpenStudy (sugarmochi):

thankyou so so much :D

OpenStudy (anonymous):

You are welcome ^_^ Feel free to ask me at any time you like :D